OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (anonymous):

multiply 3 by x and 1 multiply -2 by x and 1 3x+1=-2x-2-4 now combine like terms

OpenStudy (anonymous):

1 -1 -5 -25 and iv try the problem i cant do it

OpenStudy (anonymous):

3x+1=-2x-(-2)-4 +2x +2x 5x+1=-2-4 5x+1=-6 -1 -1 5x=5 divide by 5 on both sides x=1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :)